# Suppression of Gravitational-Wave Echoes in Diffeomorphism-Invariant Nonlocal Gravity

### v1

@@ 3 Suppression of Ringdown Echoes in Covariant Nonlocal Gravity
4 Conclusions
We have shown that gravitational–wave echo suppression is a generic and robust consequence of diffeomorphism–invariant nonlocal gravity theories constructed from analytic entire functions of the covariant d’Alembertian. The key mechanism is not model–dependent microphysics at the horizon, but rather the analyticity and Paley–Wiener bounds associated with the nonlocal regulator, which enforce exponential damping of high–frequency scattering and smear sharp structures in the effective perturbation potential.
By deriving the nonlocal master equation governing ringdown perturbations, we demonstrated that the effective photon–sphere and interior scattering amplitudes are multiplicatively filtered by the Fourier transform of the nonlocal kernel. This filtering suppresses reflection coefficients at frequencies relevant to the ringdown phase and compounds with each successive round trip between the photon sphere and the interior. As a result, any would–be echo train is exponentially attenuated with increasing frequency and echo number and is further broadened in the time domain by the intrinsic nonlocal scale.
We emphasized the distinction between regular black holes possessing true event horizons and regular horizonless compact objects. In the former case, the standard ingoing boundary condition at the horizon eliminates the inner reflection required to form a resonant cavity, independently of the ultraviolet completion. In the latter case, regularity at the center replaces the notion of a reflective surface, while analytic nonlocality strongly suppresses scattering from interior structure. In both scenarios, the two essential ingredients for observable echoes—strong inner reflectivity and a sharply defined cavity—are generically absent.
Our results have direct implications for the interpretation of current and future gravitational–wave observations. The non–detection of echoes in LIGO–Virgo–KAGRA data is fully consistent with ultraviolet–finite and quantum–gravity–motivated theories featuring regular black holes or horizonless remnants [17]. Conversely, the presence of strong, long–lived echo trains would place significant constraints on analytic nonlocal completions of gravity, favoring either non–analytic dynamics or explicit reflective boundary conditions imposed by hand.
This work highlights how fundamental properties such as diffeomorphism invariance and analyticity leave observable imprints in strong–field dynamics. Ringdown phenomenology therefore provides a complementary probe of the analytic structure and nonlocal scale of gravity, linking gravitational–wave observations to the ultraviolet completion of spacetime itself.
@@ Acknowledgments
Research at the Perimeter Institute for Theoretical Physics is supported by the Government of Canada through Industry Canada and by the Province of Ontario through the Ministry of Research and Innovation (MRI).

### v2

@@ 4 Regge–Wheeler Matching and Vanishing of the Inner Reflection Coefficient
Conclusions
We have demonstrated that diffeomorphism-invariant nonlocal gravity regulated by an analytic entire-function form factor does not produce observable gravitational-wave echoes in the merger ringdown phase of compact objects that are asymptotically Schwarzschild or Kerr. The suppression of echoes does not arise from any modification of the asymptotic quasinormal-mode spectrum, nor from a filtering of the observed gravitational-wave frequency. The ringdown frequencies measured at infinity remain those determined by the classical exterior geometry.
The essential physical mechanism is geometric. For a mode of asymptotic frequency $\omega$, the invariant proper local frequency is $\omega_{\rm loc}(r)=\frac{\omega}{\sqrt{f(r)}}$, where $f(r)=-g_{tt}(r)$. As $r\to r_{h}$ in a classical black-hole geometry, $f(r)\to 0$ and $\omega_{\rm loc}\to\infty$. Even for a horizonless but strongly redshifted compact object that mimics a black hole observationally, $f(r)$ becomes extremely small in the inner transition region, yielding $\omega_{\rm loc}\,\ell\gg 1$, where $\ell$ is the nonlocality length scale. This invariant blueshift activates the entire-function regulator in the near-horizon region. The regulator smooths out sharp inner transition layers that would otherwise act as partially reflecting boundaries. As a consequence, the inner reflection coefficient satisfies $R_{\rm in}\to 0$, eliminating the physical mechanism required for echo generation.
Echo suppression is therefore a consequence of nonlocal geometric smoothing induced by large invariant local frequencies, rather than any ad hoc frequency cutoff or damping of the observed waveform. Although the nonlocality scale is of order the Planck length, the near-horizon blueshift renders the invariant argument of the entire-function form factor large, even for low-frequency modes measured at infinity.
Within this framework, compact objects consistent with the classical exterior solutions of general relativity do not produce detectable ringdown echoes, even when the interior is rendered regular by nonlocal ultraviolet completion.
Our results highlight a conceptual distinction between local exotic compact object models and nonlocal quantum gravity. In local theories, echoes are a generic consequence of imposing hard boundary conditions near the would-be horizon. In contrast, in an analytic entire-function completion of gravity, such boundary conditions are not physically admissible. They are automatically smoothed by the same nonlocality that renders the theory ultraviolet finite and free of curvature singularities. The absence of observable echoes emerges as a prediction of nonlocal quantum gravity, rather than a fine-tuned or phenomenological assumption.
Future high-precision ringdown observations will continue to test this picture by constraining deviations from the classical quasinormal spectrum. Within the framework studied here, any confirmed detection of sharp, high-amplitude echoes would point not to a different choice of nonlocal scale, but to the breakdown of the entire-function nonlocal completion itself. In this sense, gravitational-wave echoes provide a clean observational handle on the fundamental admissibility of sharp structures in quantum gravity.
@@ Acknowledgments
I thank Ethan Thompson for helpful discussions. Research at the Perimeter Institute for Theoretical Physics is supported by the Government of Canada through Industry Canada and by the Province of Ontario through the Ministry of Research and Innovation (MRI).